# Robot Learning as an Empirical Science: Best Practices for Policy Evaluation

### v1

@@ 1 Introduction
Contribution: We propose best practices for policy evaluation to improve the science of robot learning. These include suggestions for the experimental setup, different metrics, and the analysis of the results. While we focus on evaluation with physical robots, these best practices extend well to simulation. We illustrate our points through manipulation tasks performed by physical robots.
@@ 1.1 Related Work
Experiment design and analysis in robotics and machine learning: As disciplines mature, they begin forming best practices for evaluation and reporting, to ensure further progress. Such guidelines help communicate what information is useful to the community and set expectations, which is especially useful for people entering the field. Examples for best practices in fields that are close to robot learning include a primer on conducting experiments in human-robot interaction [9], and statistical analysis for evaluation of deep reinforcement learning [10].
@@ 2.2 Initial conditions
Figure 2: Initial conditions for Example 1. Each row represents 5 evaluations with the same IC; the ICs are numbered 0-4 from the top row. The color of the text above each photo indicates the success; green is a successful rollout, red is a failure. We can see that the IC are visually consistent. Furthermore, for ICs 0,1,2 both policies always succeed, while for IC 3 they mostly fail.
@@
Figure 2(b) shows the set of initial conditions (ICs) we used to evaluate policies A and B for shirt. While the overall success rate of policy A and B are similar (72% vs 80%), we can see that the success is highly dependent on the IC; if we were not careful in controlling for IC, we might arrive at a different conclusion.
@@ 3.1 Semantic information
Capturing a notion of “success” or “failure”, these metrics are designed as Yes/No questions and include success rate (percentage of “Yes”), subgoal completion, and failures. While these metrics are task specific (even for success rate, one needs to define what success in the task means), we distinguish between metrics that are descriptive (Section 3.1.1) and metrics that are computable, for example those based on logic (Section 3.1.2).
@@ 3.1.1 Rubrics
Table 1 displays our rubric that the evaluator filled out when rolling out three different policies. Just looking at the overall task success rate makes policy C seems worthless; however, we can see that all policies were able to grasp the spatulas and flip the pancake, but policies B and C struggled with picking the pancake up and placing it on the plate.
@@
Table 1: Partial rubric for Pancake evaluation of three policies. Of note is Policy C; while the overall success rate is low (17%), it was able to complete part of the task (picking up the spatulas and flipping the pancake) 100% of the time (as denoted in bold). Looking only at overall success would deem Policy C to be an unsuccessful policy; however, that is not the full picture.
@@ 3.1.2 Signal Temporal Logic
Temporal logics [33] provide syntax and semantics for formulas that are richer than what can be captured by a predicate; for several temporal logics there exists automated evaluation procedure to assess the truth value of formulas. We propose the use of Signal Temporal Logic (STL) and both its Boolean (“Yes/No”) and quantitative semantics, also known as the robustness metric [34, 35](Section 3.2.1) as a way to create rich, computable metrics. We provide the syntax and semantics of STL in the appendix.
@@ 3.2.1 Signal Temporal Logic robustness
As mentioned above, STL has two types of semantics: Boolean and quantitative. Given an STL formula and a trajectory $X_{t}$, we can evaluate whether the formula evaluates to $True$ or $False$ (Boolean) but we can also calculate how close the formula is to satisfying or violating the formula (robustness). A positive robustness indicates that the formula is $True$, negative that it is $False$. The definition of the qualitative semantics can be found in [35]; in the following we use the RTAMT [38, 39] library to calculate the robustness of STL formulas on rollout data.
@@
We wish to evaluate whether the robot end effector is touching the table while it is pushing the bowl. To do so we look at two signals; the $z$ coordinate of the robot end effector, as calculated by the Panda using forward kinematics, and $contact$, a signal that has values of around 20 for no contact and values greater than 500 when there is contact. We create this signal from a sensor attached to the end effector. The STL specification we consider is

### v2

@@ 1 Introduction
Contribution: We propose best practices for policy evaluation to improve the science of robot learning. These include suggestions for the experimental setup, different metrics, and the analysis of the results. While we focus on evaluation with physical robots, these best practices extend well to simulation. We illustrate our points through several examples of manipulation tasks performed by physical robots.
@@ 1.1 Related Work
Experiment design and analysis in robotics and machine learning: As disciplines mature, they begin forming best practices for evaluation and reporting, to ensure further progress. Such guidelines help communicate what information is useful to the community and set expectations, which is especially useful for people entering the field. Examples for best practices in fields that are close to robot learning include a primer on conducting experiments in human-robot interaction [9], and statistical analysis for evaluation of deep reinforcement learning [10, 11].
@@ 2.2 Initial conditions
Figure 2: Initial conditions (ICs) for Example 1; Policy A on the left and Policy B on the right. Each row represents 5 evaluations with the same IC; the ICs are numbered 0-4 from the top row. The color of the text above each photo indicates the success; green is a successful rollout, red is a failure. We can see that the IC are visually consistent. Furthermore, for ICs 0,1,2 both policies always succeed, while for IC 3 they mostly fail.
@@
Figure 2(b) shows the set of initial conditions (ICs) we used to evaluate policies A and B for shirt. While the overall success rate of policy A and B are similar (72% vs 80%), we can see that the success is highly dependent on the IC; if we were not careful in controlling for IC, we might arrive at a different conclusion. For example, if we used different ICs for each policy, and by chance use mainly ICs 1,2 for one policy and ICs 3,4 for the other, we may conclude that one policy is significantly more performant than the other.
@@ 3.1 Semantic information
Capturing a notion of “success” or “failure”, these metrics are designed as Yes/No questions; they include success rate (percentage of “Yes”), subgoal completion, and failures modes. While these metrics are task specific (even for success rate, one needs to define what success in the task means), we distinguish between metrics that are descriptive (Section 3.1.1) and metrics that are computable, for example those based on logic (Section 3.1.2).
@@ 3.1.1 Rubrics
Table 1 displays our rubric that the evaluator filled out when rolling out three different policies. Just looking at the overall task success rate makes policy C seems worthless; however, we can see that all policies were able to grasp the spatulas and flip the pancake, but policies B and C struggled with picking the pancake up (last row of Table 1) and placing it on the plate (overall task success).
@@
Table 1: Partial rubric for Pancake evaluation of three policies. Of note is Policy C; while the overall success rate is low (17%), it was able to complete part of the task (picking up the spatulas and flipping the pancake) 100% of the time (as denoted in bold). Looking only at overall success would deem Policy C to be an unsuccessful policy; however, that is not the full picture, as it is able to accomplish more than half the task consistantly.
@@ 3.1.2 Signal Temporal Logic
Temporal logics [34] provide syntax and semantics for formulas that are richer than what can be captured by a predicate; for several temporal logics there exists automated evaluation procedures to assess the truth value of formulas. We propose the use of Signal Temporal Logic (STL) and both its Boolean (“Yes/No”) and quantitative semantics, also known as the robustness metric [35, 36](Section 3.2.1) as a way to create rich, computable metrics. We provide the syntax and semantics of STL in the appendix.
@@ 3.2.1 Signal Temporal Logic robustness
As mentioned above, STL has two types of semantics: Boolean and quantitative. Given an STL formula and a trajectory $X_{t}$, we can evaluate whether the trajectory evaluates to $True$ or $False$ (Boolean) but we can also calculate how close the formula is to satisfying or violating the formula (robustness). A positive robustness indicates that the formula is $True$, negative that it is $False$. Thus, STL can be used to determine both correctness and quality of policies. The definition of the qualitative semantics can be found in [36]; in the following we use the RTAMT [39, 40] library to calculate the robustness of STL formulas on rollout data.
@@
Assume we prefer the robot not touch the table when it is pushing the bowl. To automatically evaluate this property we look at two signals; the $z$ coordinate of the robot end effector, as calculated by the Panda using forward kinematics (where the $z$ axis is perpendicular to the table, $z=0$ is the table and $z$ is positive for end effector positions above the table) , and $contact$, a signal that has values of around 20 for no contact and values greater than 500 when there is contact. We create this signal from a sensor attached to the end effector. The STL specification capturing the required property is